### v1

# Recovering Hardy spaces from optimal domains of integration operators

Abstract.
We study the optimal domains for bounded Volterra integration operators $T_{g}$ between distinct Hardy spaces $H^{p}$ and $H^{q}$ of the unit ball. It is shown that the intersection of the optimal domains is equal to $H^{p}$ if $p>q$, whereas if $p<q$, we show that this intersection is genuinely larger. In the unit disk, this problem was recently solved for $p=q$ by Bellavita, Daskalogiannis, Nikolaidis and Stylogiannis.
Key words and phrases:
Volterra operator, Hardy space, optimal domain
2020 Mathematics Subject Classification:
Primary 47B35; Secondary 30H05, 46E20.
S. Eskandari was funded by the postdoctoral scholarship JCK22-0052 from the Kempe Foundations.
1. Introduction
Given $z=(z_{1},z_{2},...,z_{n})$ and $\zeta=(\zeta_{1},\zeta_{2},...,\zeta_{n})$, two vectors of the complex space $\mathbb{C}^{n}$, we denote
$\langle z,\zeta\rangle=\sum_{k=1}^{n}z_{k}\overline{\zeta_{k}}.$
We set $|z|=\sqrt{\langle z,z\rangle}$ and let $\mathbb{B}_{n}=\{z\in\mathbb{C}^{n}:|z|<1\}$ be the open unit ball, and $\mathbb{S}_{n}=\partial\mathbb{B}_{n}=\{z\in\mathbb{C}^{n}:|z|=1\}$ its boundary, the unit sphere. Let $H(\mathbb{B}_{n})$ denote the space of all holomorphic functions on $\mathbb{B}_{n}$, equipped with the topology of uniform convergence on compact sets. We use notation $R$ for the radial derivative:
$Rf(z)=\sum_{k=1}^{n}z_{k}\partial_{z_{k}}f(z).$
Given $g\in H(\mathbb{B}_{n})$ the operators $T_{g}$ and $S_{g}$ acting on $H(\mathbb{B}_{n})$ given by
$T_{g}f(z)=\int_{0}^{1}f(tz)Rg(tz)\frac{dt}{t},\quad S_{g}f(z)=\int_{0}^{1}Rf(tz)g(tz)\frac{dt}{t}$
are well-defined. The operator $T_{g}$ is called the (generalized) Volterra integration operator, while $S_{g}$ is called the Volterra companion operator. Pommerenke [19] was the first to study $T_{g}$ in one dimension; the variant we use for the unit ball appears first in the work of Hu [15].
Assume now that $X$ and $Y$ are two topological vector spaces of holomorphic functions on $\mathbb{B}_{n}$. Inspired by [21], we consider
[…]
Clearly $V$ is always a vector space. If $X$ and $Y$ are quasi-Banach spaces, then $V$ is a quasinormed space equipped with the topology coming from $\|g\|_{V}=\|T_{g}\|_{X\to Y}+|g(0)|$. Moreover, if convergence in both $X$ and $Y$ implies uniform convergence on compact sets (that is, $X$ and $Y$ continuously embed into $H(\mathbb{B}_{n})$), and they both contain the constants, then $V$ is also complete. This fact is proven in [21] in the setting of one complex variable when $T_{g}$ acts $X\to X$, with a slightly less restrictive requirements on $X$. The reader should have no difficulties verifying that it also holds for $T_{g}:X\to Y$ under the assumptions above.
The classical problem for $T_{g}$ is to study its boundedness between two Hardy spaces. We recall that for $0<p<\infty$, the Hardy space $H^{p}$ consists of $f\in H(\mathbb{B}_{n})$ with
$\|f\|_{H^{p}}=\sup_{0\leq s<1}\left(\int_{\mathbb{S}_{n}}|f(s\zeta)|^{p}d\sigma(\zeta)\right)^{1/p}<\infty,$
where $d\sigma$ is the normalized surface measure on $\mathbb{S}_{n}$. When $p\geq 1$, the norm $\|\cdot\|_{H^{p}}$ makes $H^{p}$ a Banach space. When $0<p<1$, the space $H^{p}$ is a complete quasi-Banach space. Moreover, in that case $d_{p}(f,g)=\|f-g\|_{H^{p}}^{p}$ is a complete translation invariant metric on $H^{p}$, making it an F-space in the sense it is used by Rudin [20] (they are not locally convex). For basic properties of these spaces, we refer the reader to [13, 23, 24].
For $T_{g}:H^{p}\to H^{q}$, the spaces $V$ are completely known. In fact (see and [18], or [4, 5] for earlier proofs in the case $n=1$),
  • $p=q$: $V=BMOA$;
  • $p<q$: $V=\mathcal{B}^{\alpha}$;
  • $p>q$: $V=H^{r}$.
Here r=pq/(pq)r=pq/(p-q), $\alpha=1+n/q-n/p$ and $\mathcal{B}^{\alpha}$ is the Bloch type space consisting of analytic functions $g$ with
$\|g\|_{\mathcal{B}^{\alpha}}=\sup_{z\in\mathbb{B}_{n}}(1-|z|^{2})^{\alpha}|Rg(z)|<\infty,$
and we understand this space to consist of constants only, if $\alpha<0$ (in which case $T_{g}$ is the zero operator). The space $BMOA$ is the space of holomorphic functions of bounded mean oscillation – the precise definition is not used in this paper, but it can be found in [23] and in [13, 24] for $n=1$.
Given $X$, $Y$, and $g\in V$, we may ask which holomorphic functions $f$ are mapped into $Y$ by $T_{g}$. This vector space of holomorphic functions will be called the optimal domain of $T_{g}$ and denoted
$g:Y]=\{f\in H:T_{g}f\in Y\}.$
This space $[g:Y]$ can be equipped with the quasinorm $\|f\|_{[g:Y]}=\|T_{g}f\|_{Y}$. The study of optimal domains for integration operators was initiated by Curbera and Ricker [10, 11], who looked into this concept for the Cesáro operator. In the recent years, there has been a number of interesting papers on the topic, see [2, 3, 7], or [6] for meromorphic optimal domains. We mention that all the aforementioned papers are written in the setting of the unit disk. We also remark that our notation differs slightly from them. However, in the context of Volterra operators only, this notation should not confuse the reader. We have chosen it because it makes the notation $[V:Y]$ that is given below more convenient.
From the definitions it follows that we always have $X\subseteq[g:Y]$ for $g\in V$, and if, for instance $X=Y=H^{p}$ and $n=1$, the inclusion is known to be proper [7]. However, another interesting question is whether the intersection of all the optimal domains agrees with $X$. That is, whether $[V:Y]=X$, where
$V:Y]=\bigcap_{g\in V}[g:Y].$
We understand this problem in the following way. While $X$ and $Y$ uniquely define $V$, when is it true that $Y$ and $V$ uniquely define $X$? Or, in other words, can $X$ be recovered from the optimal domains of bounded Volterra operators $X\to Y$.
Before going further, let us offer a simple example. Let $X=A(\mathbb{B_{n}})$ be the ball algebra and $Y=\mathcal{B}^{1}$ be the standard Bloch space. It is easy to see that $V=\mathcal{B}^{1}$, and that for every $g\in V$ one has $H^{\infty}\subseteq[g:\mathcal{B}^{1}]$, so clearly $X\subsetneq[V:Y]$ in that case. But one might argue that an example such as this one is tailor-made for the identity to fail. So what happens in the Hardy spaces?
In [7] it is proven that
$BMOA:H^{p}]=H^{p},\quad p\geq 1,$
when $n=1$. In this work, we work in the setting of the unit ball and prove that if $p>q$ and r=pq/(pq)r=pq/(p-q), then
$H^{r}:H^{q}]=H^{p},$
which means that $H^{p}$ can be recovered from the optimal domains of bounded $T_{g}:H^{p}\to H^{q}$ for $p>q$.
On the other hand, when $p<q$ and $\alpha=1+n/q-n/p$, we prove
$H^{p}\subsetneq[\mathcal{B}^{\alpha},H^{q}].$
This means that $H^{p}$ cannot be recovered from optimal domains of bounded $T_{g}:H^{p}\to H^{q}$ for $p>q$. Other properties of optimal domains of this type were very recently considered in [3] for $n=1$.
The proofs of these theorems are presented in the next section. We will also study the boundedness of $S_{g}$ between two Hardy spaces, as this is key to studying optimal domains using the approach in [7]. We also discuss some further generalizations and open problems.
As one might expect, we will need comparisons between different quantities. As is standard, by $A\lesssim B$ (or equivalently $B\gtrsim A$), we mean that there exists $C>0$ so that $A\leq CB$. If both $A\lesssim B$ and $B\gtrsim A$ are true, we write $A\asymp B$.
2. Main results
We will go straight into the proofs of the results mentioned in the Introduction.
Theorem 1.
Let $1<q<p<\infty$ and r=pq/(pq)r=pq/(p-q). Then $H^{p}=[H^{r}:H^{q}]$.
Proof.
We follow the reasoning in [7]. Let $g\in V(H^{p},H^{q})=H^{r}$, and $f\in[H^{r}:H^{q}]$. Now, $T_{g}f$ belongs to $H^{q}$, or equivalently, $S_{f}g\in H^{q}$. This means by the closed graph theorem (for F-spaces if necessary; see [20]) that $S_{f}:H^{r}\to H^{q}$ is bounded. The claim now follows from the next theorem. ∎
We will formulate the following theorem in the standard form. That is, for $S_{g}:H^{p}\to H^{q}$. Note that if $p>q$, the roles of $p$ and $r$ are symmetric, so we will be able to finish the proof once this result is obtained.
Theorem 2.
Let $1<q<p<\infty$, r=pq/(pq)r=pq/(p-q) and $g\in H(\mathbb{B}_{n})$. Then $S_{g}:H^{p}\to H^{q}$ is bounded if and only if $g\in H^{r}$. Moreover, if $g(0)=0$, we have
$\|T_{g}\|_{H^{p}\to H^{q}}\asymp\|g\|_{H^{r}}.$
Proof.
We note that this result is probably known to experts, at least when $n=1$. Since we have been unable to find its proof in the literature, one is given here.
We first consider the sufficiency. In fact, it follows easily from corresponding results for $T_{g}$ and the multiplication operator $M_{g}$. We provide a direct proof, similar to one for $T_{g}$ in [18], by methods that will be used for the necessity. For $\gamma>1$ and $|\zeta|=1$, we denote by $\Gamma(\zeta)=\Gamma_{\gamma}(\zeta)$ the approach region
$\Gamma_{\gamma}(\zeta)=\left\{z\in\mathbb{B}_{n}:|1-\langle z,\zeta\rangle|<\frac{\gamma}{2}(1-|z|^{2})\right\}.$
For $\beta\in\mathbb{R}$ we write $dV_{\beta}(z)=(1-|z|^{2})^{\beta}dV(z)$, where $V$ is the normalized $2n$-dimensional Lebesgue measure on $\mathbb{B}_{n}$. Now assume $g\in H^{r}$ with r=pq/(pq)r=pq/(p-q). If $f\in H^{p}$, then by the area function description of Hardy spaces [1, 18] we have
$\|S_{g}f\|^{q}_{H^{q}}$ $\asymp\int_{\mathbb{S}_{n}}\left(\int_{\Gamma(\zeta)}|Rf(z)g(z)|^{2}dV_{1-n}(z)\right)^{q/2}d\sigma(\zeta)$
$\lesssim\int_{\mathbb{S}_{n}}\left(\sup_{z\in\Gamma(\zeta)}|g(z)|\right)^{q}\left(\int_{\Gamma(\zeta)}|Rf(z)|^{2}dV_{1-n}(z)\right)^{q/2}d\sigma(\zeta).$
We use Hölder’s inequality and get
$\|S_{g}f\|^{q}_{H^{q}}\lesssim$ $\left(\int_{\mathbb{S}_{n}}\left(\sup_{z\in\Gamma(\zeta)}|g(z)|\right)^{pq/(p-q)}d\sigma(\zeta)\right)^{1-q/p}$
$\times\left(\int_{\mathbb{S_{n}}}\left(\int_{\Gamma(\zeta)}|Rf(z)|^{2}dV_{1-n}(z)\right)^{p/2}d\sigma(\zeta)\right)^{q/p}.$
Equivalently, this means
$\|S_{g}f\|^{q}_{H^{q}}\lesssim\|g\|^{q}_{H^{r}}\|f\|^{q}_{H^{p}},$
giving the sufficiency.
For the necessity, we proceed with the argument from [17], based on the Kahane and Khinchine inequalities [16], atomic decomposition [18], area function description of Hardy spaces [1, 18], and factorization of tent spaces [8, 9, 17]. The proof follows the outline of either Theorem 7 or Theorem 8 of [17], except that it is simpler. The proof is sketchy in the sense that we do not discuss the preliminaries on tent spaces is this paper; an interested reader can take a look at [17].
Let $Z=\{a_{k}\}\subseteq\mathbb{B}_{n}$ be a $\rho$-lattice in the Bergman metric. The most of the argument below works for any $\rho>0$, but in the very last stage we will assume that $\rho$ is very small. Let $b>n\max(1,2/p)$. We consider a family of test functions of the form
$f_{t}(z)=\sum_{k}\lambda_{k}r_{k}(t)\left(\frac{1-|a_{k}|^{2}}{1-\langle z,a_{k}\rangle}\right)^{b},$
where the sequence $\{\lambda_{k}\}$ belongs to $T_{2}^{p}$, a tent space of sequences indexed by the lattice, and $r_{k}:[0,1]\to\{-1,+1\}$ are the Rademacher functions. By the area function description of the Hardy spaces, integrating with respect to t, where $t\in[0,1],$ and using Fubini’s theorem, we have
$\int_{\mathbb{S}_{n}}\int_{0}^{1}$ $\left(\int_{\Gamma(\zeta)}\left|g(z)\sum_{k}\lambda_{k}r_{k}(t)|\langle z,a_{k}\rangle|\dfrac{(1-|a_{k}|^{2})^{b}}{(1-\langle a_{k},z\rangle)^{b+1}}\right|^{2}dV_{1-n}(z)\right)^{q/2}dt\,d\sigma(\zeta)$
$\lesssim\|S_{g}\|^{q}\,\|\lambda\|^{q}_{T^{p}_{2}}.$
Now, using Kahane’s inequality and Fubini’s theorem this gives
$\int_{\mathbb{S}_{n}}$ $\left(\int_{\Gamma(\zeta)}\int_{0}^{1}\left|g(z)\langle z,a_{k}\rangle\sum_{k}\lambda_{k}r_{k}(t)\dfrac{(1-|a_{k}|^{2})^{b}}{(1-\langle z,a_{k}\rangle)^{b+1}}\right|^{2}dt\,dV_{1-n}(z)\right)^{q/2}\,d\sigma(\zeta)$
$\lesssim\|S_{g}\|^{q}\,\|\lambda\|^{q}_{T^{p}_{2}}.$
We use Khinchine’s inequality to obtain
$\int_{\mathbb{S}_{n}}$ $\left(\int_{\Gamma(\zeta)}\sum_{k}|\lambda_{k}|^{2}|\langle z,a_{k}\rangle|^{2}|g(z)|^{2}\dfrac{(1-|a_{k}|^{2})^{2b}}{|1-\langle z,a_{k}\rangle|^{2b+2}}\,dV_{1-n}(z)\right)^{q/2}\,d\sigma(\zeta)$
$\lesssim\|S_{g}\|^{q}\,\|\lambda\|^{q}_{T^{p}_{2}}.$
Replacing the integral over $\Gamma(\zeta)$ by an integral over Bergman metric balls $B(a_{k},\rho)$, and switching to an approach region with a smaller aperture, also denoted by $\Gamma(\zeta)$,
$\int_{\mathbb{S}_{n}}$ $\left(\sum_{{a_{k}}\in\Gamma(\zeta)}|\lambda_{k}|^{2}\int_{B(a_{k},r)}|\langle z,a_{k}\rangle|^{2}|g(z)|^{2}\dfrac{(1-|a_{k}|^{2})^{2b}}{|1-\langle z,a_{k}\rangle|^{2b+2}}\,dV_{1-n}(z)\right)^{q/2}\,d\sigma(\zeta)$
$\lesssim\|S_{g}\|^{q}\,\|\lambda\|^{q}_{T^{p}_{2}}.$
We proceed by using subharmonicity and $|1-\langle z,a_{k}\rangle|=1-|a_{k}|^{2}$ for every $z\in D(a_{k},r)$, arriving at the estimate
(2.1) $\int_{\mathbb{S}_{n}}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}|\lambda_{k}|^{2}|a_{k}|^{2}|g(a_{k})|^{2}\right)^{q/2}\,d\sigma(\zeta)\lesssim\|S_{g}\|^{q}\,\|\lambda\|^{q}_{T^{p}_{2}}.$
Denoting $\nu_{k}=|a_{k}|^{2}|g(a_{k})|,$ we want to prove that $\{\nu_{k}\}\in T^{r}_{\infty}$. This would lead to a discrete way to express that $g\in H^{r}$.
Let $s>0$ be big enough so that $2s>1$ and $qs>1$ hold. We are done if we can prove that $\{\nu_{k}^{1/s}\}\in T^{rs}_{\infty}$. By duality of tent spaces, since $rs>1$, this can be shown if we prove
$\left|\sum_{k}\mu_{k}\nu_{k}^{1/s}(1-|a_{k}|^{2})^{n}\right|\lesssim\|\mu\|_{T^{(rs)^{\prime}}_{1}}$.
To this end, we assume that $\lambda_{k}$ and $\tau_{k}$ are positive. Indeed, since $\nu_{k}$ are positive, the above dual pairing is clearly at its biggest for a positive sequence $\{\mu_{k}\}$, and the sign change on individual elements does not affect the norm of the corresponding tent space. We get
$\sum_{k}\tau_{k}\lambda_{k}^{1/s}\nu_{k}^{1/s}(1-|a_{k}|^{2})^{n}\asymp\int_{\mathbb{S}_{n}}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}\tau_{k}\lambda_{k}^{1/s}\nu_{k}^{1/s}\right)\,d\sigma(\zeta).$
Next, using Hölder’s inequality, yields
$\sum_{k}\tau_{k}\lambda_{k}^{1/s}\nu_{k}^{1/s}(1-|a_{k}|^{2})^{n}$
$\lesssim$ $\int_{\mathbb{S}_{n}}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}\tau_{k}^{2s/(2s-1)}\right)^{(2s-1)/2s}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}\lambda_{k}^{2}\nu_{k}^{2}\right)^{1/(2s)}\,d\sigma(\zeta).$
Consider $2s/(2s-1)=(2s)^{\prime}$ and doing another Hölder’s inequality for $qs$ and $qs/(qs-1)=(qs)^{\prime}$ gives us now that the expression is dominated by
$\left(\int_{\mathbb{S}_{n}}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}\tau_{k}^{\frac{2s}{2s-1}}\right)^{\frac{(qs)^{\prime}}{(2s)^{\prime}}}\,d\sigma(\zeta)\right)^{\frac{1}{(qs)^{\prime}}}\left(\int_{\mathbb{S}_{n}}\left(\sum_{{a_{k}}\in\Gamma(\zeta)}\lambda_{k}^{2}\nu_{k}^{2}\right)^{\frac{q}{2}}\,d\sigma(\zeta)\right)^{\frac{1}{qs}}.$
The first part is the norm $\|\tau\|_{T^{(qs)^{\prime}}_{(2s)^{\prime}}(Z)}.$ By using 2.1 for the second part, to get the upper bound $\|S_{g}\|^{1/s}\|\lambda\|^{1/s}_{T^{p}_{2}}=\|S_{g}\|^{1/s}\|\lambda\|_{T^{ps}_{2s}}$. Altogether we have
$\sum_{k}\tau_{k}\lambda_{k}^{1/s}\nu_{k}^{1/s}(1-|a_{k}|^{2})^{n}\lesssim\|S_{g}\|^{1/s}\|\lambda^{1/s}\|_{T^{ps}_{2s}}\|\tau\|_{T^{(qs)^{\prime}}_{(2s)^{\prime}}}.$
By factorization of tent spaces (essentially due to Cohn and Verbitsky [8], but the present variant is taken from [17]) and considering $\mu_{k}=\tau_{k}\lambda_{k}^{1/s}$ as an arbitrary positive element of the space $T^{(rs)^{\prime}}_{1}$. By duality we have $\{\nu_{k}^{1/s}\}\in T^{rs}_{\infty}$. By the discussion and calculation above, this leads to $\|\nu\|_{T^{r}_{\infty}}\lesssim\|S_{g}\|,$ which is what we wanted to prove.
Now, choosing $\rho>0$ small enough and invoking Lemma 3 of [17], we get
$\int_{\mathbb{S}_{n}}\left(\sup_{z\in\Gamma(\zeta)}|z_{i}|^{2}|g(z)|\right)^{r}d\sigma(\zeta)<\infty$
for every $i\in\{1,2,...,n\}$ (since $z_{i}^{2}g(z)$ is holomorphic), yielding that
(2.2) $\int_{\mathbb{S}_{n}}\left(\sup_{z\in\Gamma(\zeta)}|z|^{2}|g(z)|\right)^{r}d\sigma(\zeta)<\infty.$
By the admissible maximal function description of Hardy spaces, we see that $H^{r}$ and the tent space of holomorphic functions defined by (2.2) are in bijective correspondence. By the Bounded Inverse Theorem for F-spaces, we get that $g\in H^{r}$ with its norm dominated by the operator norm of $S_{g}$.
To present the next theorem, we will need to introduce the standard weighted Bergman spaces. Let $\beta>-1$ and $0<p<\infty$. The weighted Bergman space $A^{p}_{\beta}$ consists of $f\in H(\mathbb{B}_{n})$ with
$\|f\|_{A^{p}_{\beta}}=\left(\int_{\mathbb{B}_{n}}|f(z)|^{p}dV_{\beta}(z)\right)^{1/p}<\infty.$
See [23, 24] for some basic theory of $A^{p}_{\beta}$. We will need the following inclusion between Hardy and Bergman spaces for $p<s$.
(2.3) $H^{p}\subsetneq A^{s}_{n\frac{s}{p}-n-1}.$
In the unit disk this result goes back at least to 1932 paper of Hardy and Littlewood [14]. It is also a special case of Duren’s theorem [12]. It is moreover known that the spaces $A^{s}_{n\frac{s}{p}-n-1}$ increase in size as $s$ increases. A simple proof of (2.3) can be found in [23] (see Theorem 4.48).
Theorem 3.
Let $1<p<q<\infty$ and $\alpha=1+n/q-n/p$. Then one has $H^{p}\subsetneq[\mathcal{B}^{\alpha}:H^{q}]$.
Proof.
Assume $p<q$ and $g\in\mathcal{B}^{\alpha}$. We consider two cases, $q>2$ and $q\leq 2$. For the case $q>2$, let $s\in(2,q)\cap(p,q)$. Then we have $2<s<q$ and from Theorem 1 (1) of [17], it follows that $T_{g}:A^{s}_{n\frac{s}{p}-n-1}\to H^{q}$ is bounded. Therefore $A^{s}_{n\frac{s}{p}-n-1}\subseteq[g:H^{q}]$.
Let us next assume that $q\leq 2$. Here we could also proceed by taking $s\in(p,q)$, in which case $s\leq\min\{2,q\}$. But in fact we can let $s=q$, obtaining a larger space. According to case (1) of Theorem 1 in [17] (in fact, in the unit disk this case was obtained earlier by Wu [22]), the boundedness of $T_{g}:A^{q}_{n\frac{q}{p}-n-1}\to H^{q}$ is equivalent to $g\in\mathcal{B}^{\alpha}$.
Now, by using (2.3), the space $H^{p}$ embeds into $A^{s}_{n\frac{s}{p}-n-1}$ or $A^{q}_{n\frac{q}{p}-n-1}$, whichever we want. The inclusion is also strict as Bergman functions need not have boundary values almost everywhere. Therefore the proof is complete. ∎
Remark 4.
While it is interesting that $H^{p}$ cannot be recovered from the optimal domains of bounded Volterra operators in this case, the result is not too surprising. Indeed, if $p<q$ and $\alpha=1+n/q-n/p<0$, then $V$ consists of constant functions. In this case $V:H^{q}]=H(\mathbb{B}_{n})$, which is clearly much bigger than $H^{p}$, and the initial domain space of $T_{g}$ can obviously not be recovered as intersection of the optimal domains.
Let $w\geq 0$ be a measurable function. We denote by $AT^{p}_{q}(w)$ the weighted tent space (originally introduced by Coifman, Meyer and Stein [9]) consisting of analytic functions $f$ with
$\|f\|_{AT^{p}_{q}(w)}=\left(\int_{\mathbb{S}_{n}}\left(\int_{\Gamma(\zeta)}|f(z)|^{2}w(z)dV_{1-n}(z)\right)^{p/q}d\sigma(\zeta)\right)^{1/p}.$
Given an analytic $g$, using the area description of $H^{q}$, one obtains
$g:H^{q}]=AT^{q}_{2}(|Rg|^{2}).$
The findings of this paper can be expressed in the following form.
Corollary 5.
Let $0<p,q<\infty$. The following are true:
  • i.
    If $p>q$ and r=pq/(pq)r=pq/(p-q), then
    $H^{p}=\bigcap_{g\in H^{r}}AT^{q}_{2}(|Rg|^{2});$
  • ii.
    If $p<q$ and $\alpha=1+n/q-n/p$, then
    $H^{p}\subsetneq\bigcap_{g\in\mathcal{B}^{\alpha}}AT^{q}_{2}(|Rg|^{2}).$
As mentioned in the beginning of the paper, when $n=1$ the case $p=q\geq 1$ was solved in [7]. However, we have not been able to show that $H^{p}=[BMOA:H^{p}]$ when $n\geq 2$, although it is true for $n=1$.